OpenStudy (anonymous):

use the implicit differentiation to find dy/dx of 4x^2-8xy=y^2-8x

OpenStudy (shamil98):

Use the power and product rules :)

OpenStudy (shamil98):

Do you know how to do implicit diff tho?

OpenStudy (anonymous):

nope

OpenStudy (shamil98):

Differentiate with respect to x, what do you get? you are supposed to find y' \[\huge 4x^2-8xy=y^2-8x\] \[\huge 8x - 8y -8xy' = 2yy' -9\] Group the terms with y' and solve for y'

OpenStudy (shamil98):

that should be - 8* at the end

OpenStudy (shamil98):

\[\huge 8x - 8y + 8 = y' (8x + 2y)\] \[\huge \frac{ 8(x-y+1) }{ 8x + 2y } = y'\] you understand?..